can somebody help me solve this problem and show me the work so i understand how to do it
5-2(x+6)=1

Answer by Edwin McCravy(20059)

5-2(x+6)=1

We copy down the 5 under that

5-2(x+6)=1
5

Now we multiply the -2 by the x and get -2x and write that down
after the 5:

5-2(x+6)=1
5-2x

Now we multiply the -2 by the +6 and get -12 and write that down
after the 5-2x:

5-2(x+6) = 1
5-2x-12

Now we bring down the equal sign and the 1:

5-2(x+6) = 1
5-2x-12  = 1

Now on the next line we combine the 5 and the -12 getting -7 and 
write that on the next line without the 5 and the -12, but we 
bring down the -2x, the = sign and the 1 and we have

5-2(x+6) = 1
5-2x-12  = 1
  -7-2x  = 1

Next we add 7 to both sides canceling the -7 on the left and giving
8 on the right, and we bring down only the -2x, the = and the 1 to the 
next line:

5-2(x+6) = 1
5-2x-12  = 1
  -7-2x  = 1
    -2x  = 8

Finally we divide both sides by -2, getting x on the left and -4 on
the right, and we put that on the bottom line.

5-2(x+6) = 1
5-2x-12  = 1
  -7-2x  = 1
    -2x  = 8
      x  = -4 

Edwin